A tuna producer claims that its cans of tuna contain a standard amount of tuna. It has been determined that of 200 cans manufactured, 92% contain the standard amount. An inspector will only accept a shipment if at least 175 out of 200 sample cans contain more than the standard amount. What is the probability that the shipment is accepted?


1
Expert's answer
2022-01-24T16:23:19-0500

Let "X=" the number of cans that contain the standard amount of tuna: "X\sim Bin (n, p)."

Given "n=200, p=0.92."

If the binomial probability histogram is not too skewed, "X" has approximately a normal distribution with "\mu=np" and "\sigma=\sqrt{npq}."

In practice, the approximation is adequate provided that both "np\geq 10" and "nq\geq 10," since there is then enough symmetry in the underlying binomial

distribution.

"np=200(0.92)=184\geq 10"

"nq=n(1-p)=200(1-0.92)=16\geq 10"

Hence we can use Normal approximation to the Binomial: "X\sim N(\mu, \sigma^2)"


"\mu=np=200(0.92)=184"

"\sigma^2=npq=200(0.92)(1-0.92)=14.72"

Use the continuity correction factor


"P(X\geq 175)=P(X>175-0.5)"

"=1-P(X\leq174.5)=1-P(Z\leq\dfrac{174.5-184}{\sqrt{14.72}})"

"\approx1-P(Z\leq-2.4761)\approx0.9934"

The probability that the shipment is accepted is 0.9934.